12) A heat engine operates between a source at 550°C and a sink at 25°C. If heat is supplied

tothe heat engine at a steady rate of 1200 kJ/min, determine the maximum power output

of this heat engine.


1
Expert's answer
2022-02-09T05:13:01-0500

"\frac{W}{Q}=1-\frac{T_2}{T_1}"

"W= (1-\frac{298}{823})\times 1200 =765.49""\frac{KJ}{min}" =12.75 kW